What is relation b/w mole fraction of solute and molality of a solution?

Grade:11

1 Answers

Sommay shivangi
6 years ago
Mass of solution =na.Ma+ nb.MbVolume= mass/ density= (na.Ma+nb.Mb)/dMolarity=(no. Of moles of solute*1000)/volume of solutionM= (nb*1000)/(na.Ma+nb.Mb)/dM= (nb*1000*d)/(na.Ma+nb.Mb)Div. By na+nb=> M=(xb*1000*d)/(xa.Ma+xb.Mb)Where, na- no.of moles of solventnb- no. Of moles of soluteMa- molecular mass of solventMb-molecular mass of soluted- densityxa- mole fraction of solventxb- mole fraction of soluteM- molarity of solution